OpenStudy (anonymous):

Find the product. (7-10x)^2

OpenStudy (anonymous):

@Vlery

OpenStudy (vlery):

find x?

OpenStudy (anonymous):

I guess it wants you to find the product of the whole equation do you want me to post answer choices?

OpenStudy (vlery):

yes please

OpenStudy (anonymous):

ok hold on just one sec

OpenStudy (anonymous):

B. 49-149x+100x C. 49+100x D. 49-140x-100x

OpenStudy (vlery):

wheres a? lol

OpenStudy (anonymous):

it was wrong lol

OpenStudy (anonymous):

i think it is between B or D

OpenStudy (anonymous):

@Vlery are you there?

OpenStudy (vlery):

yea I'm here

OpenStudy (anonymous):

I am thinking B what do you think

OpenStudy (vlery):

I'm pretty sure it's C

OpenStudy (anonymous):

i guess i will try C. and see i thought it was B.

OpenStudy (anonymous):

opps hold on a sec

OpenStudy (anonymous):

B. 49-140x+100x^2 C. 49 + 100x^2 D. 49- 140x-100x^2

OpenStudy (anonymous):

i saw where i messed up on the answers

OpenStudy (uri):

You use this formula: (a-b) ^2 =a^2-2ab +b^2 (7-10x)^2 =49-140+100x^2

OpenStudy (anonymous):

so it is B. @uri

OpenStudy (uri):

Ye :)

OpenStudy (anonymous):

Thank you I thought so :)

OpenStudy (anonymous):

Thank you @Vlery also :)

OpenStudy (vlery):

sorry I couldn't help much

OpenStudy (anonymous):

NO its fine would you like to help me with another?